OpenStudy (anonymous):

I need to check my answer please! I chose B. Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

yeah generate the first 5 terms... where exactly are you stuck ?

ganeshie8 (ganeshie8):

Notice that you're already given the first two terms : ``` f(1) = 0 f(2) = 1 ```

OpenStudy (anonymous):

im not stuck i just want to make sure im correct

ganeshie8 (ganeshie8):

So the sequence stars like below : 0, 1, ...

OpenStudy (anonymous):

and i know it's not A

ganeshie8 (ganeshie8):

thats a very good start !

ganeshie8 (ganeshie8):

lets generate the third term

ganeshie8 (ganeshie8):

looks we're given a formula for generating terms ?

ganeshie8 (ganeshie8):

` f(n) = f(n - 1) + f(n - 2)` plugin n = 3, you wil get the 3rd term : f(3) = f(3-1) + f(3-2) = f(2) + f(1) = ?

OpenStudy (anonymous):

the thing is that i dont really know how to add f(2) + f(1) because the lesson did not really explain everything to me... i read it over again like 3 times.. :(

ganeshie8 (ganeshie8):

f(2) = second term f(1) = first term

ganeshie8 (ganeshie8):

just add the first and second terms to get the third term :)

OpenStudy (anonymous):

1

ganeshie8 (ganeshie8):

Perfect ! f(3) = third term = 1

ganeshie8 (ganeshie8):

So far the sequence is : 0, 1, 1,...

ganeshie8 (ganeshie8):

see if u can generate the 4th term

OpenStudy (anonymous):

2

ganeshie8 (ganeshie8):

Yes ! 0, 1, 1, 2,...

ganeshie8 (ganeshie8):

what about 5th term ?

OpenStudy (anonymous):

3

ganeshie8 (ganeshie8):

Yep ! good job !!

OpenStudy (anonymous):

thanks! :)